1. Part A What is the resistance of a 1.5-m-long copper wire that is 0.30 mm in diameter? Submit My Answers Give Up Part B What is the resistance of a 10-cm-long piece of iron with a 1.5 mm x 1.5 mm square cross section? Express your answer using two significant figures Submit My Answers Give Up

2.

3.
Answer #1

R = rho*L/A

rho = restivity of copper = 1.68 x 10^-8 ohm-m

L = lenght

A = area

part a)

A = pi*r^2

r = 0.15 mm = 0.15 x 10^-3 m

L = 1.5 m

R = 0.3565 ohm

part b )

L = 10 cm = 0.1 m

A = 1.5 x 10^-3 * 1.5 x 10^-3

R = 7.5 x 10^-4 ohm

part 1 )

R = rho*L/A

now L' = 2L

r' = 2r

R' = rho*L'/pi*r'^2

R' = rho*2L/pi*4r^2

R' = R/2